OpenStudy (anonymous):

Medal+Fanning~ Your piggy bank has a total of 47 coins in it; some are dimes and some are nickels. If you have a total of $3.95, how many nickels do you have? Numerical Answers Expected!

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

I think this one is 15 nickels

OpenStudy (anonymous):

n + d = 47 --- > d = 47 - n .05n + .10d = 3.95 sub 47 - n in for d in the 2nd equation .05n + .10(47 - n) = 3.95 .05n + 4.70 - .10n = 3.95 -.05n = 3.95 - 4.70 -.05n = -.75 n = -.75/-.05 n = 15 you are correct...there are 15 nickels

OpenStudy (anonymous):

@kelliegirl33 for the numeric answer would it be written as n=15 or just 15?